Circular No. 7382

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA SCUTI 2000 Katsumi Haseda, Aichi, Japan, reports his discovery of an apparent

nova (mag 10.6) on unfiltered Kodak T-Max 400 films taken on Mar. 5.810 UT with 0.10-m f/4 twin patrol cameras. Exposures by Haseda on Mar. 13.82 give mag 11.6, and an earlier image on Feb. 6.844 shows the star at mag 12.0. An unfiltered CCD image taken by M. Uemura and T. Kato, Kyoto University, on Mar. 14.85 yields mag 11.2 and the following position for N Sct 2000 (measured by H. Yamaoka, Kyushu University): α = 18^{h}34^{m}03^{s}.16, δ = −14^{o}45′11′′.5 (equinox 2000.0). Yamaoka further reports that spectra of the star taken on Mar. 16.81 by M. Fujii, Kurashiki, Okayama, with a 0.28-m reflector show strong Hα emission (FWHM = 940 km/s, steeper blueward), indicating that this is indeed a nova; emission lines of Hβ and Fe II are also detected with blueward steepness or weak absorption. An image obtained by J. Quinn and P. Garnavich with the 1.8-m Vatican Advanced Technology Telescope at Mt. Graham on Mar. 16.5 shows the nova at R = 11.3 ± 0.1, and they note that it appears to be quite blue from UBVRI images; Garnavich adds that comparison with a IIIa-F (red) U.K. Schmidt plate taken on 1988 Apr. 23 shows a clumping of very faint stars near the nova’s position, but that if a precursor is present, it must be fainter than mag 18. Additional magnitudes (forwarded by both Haseda and Yamaoka) from T-Max 400 films taken with 0.10-m twin patrol cameras by K. Takamizawa, Saku-machi, Nagano: Feb. 9.84, 12.2; 11.84, 12.0; 16.83, 11.8; Mar. 2.80, 9.9; 8.78, 12.0. Numerous other images by Takamizawa and Haseda from 1998 Aug. 16 to 1999 Nov. 10.38 show nothing at the location of N Sct 2000 (limiting mag 13.0–14.5).

SUPERNOVA 2000H IN IC 454 K. Krisciunas and A. Rest, University of Washington, report prelimi-

nary photometry of SN 2000H obtained with the 3.5-m telescope at Apache Point Observatory (photometric uncertainties estimated as ± 0.03 mag): Feb. 10.34 UT, V = 17.30, B − V = +0.59; 14.09, V = 17.40, B − V = +0.77, V − R = +0.43; Mar. 11.11, 18.58, +1.28, +0.67.

KL DRACONIS N. N. Samus, Sternberg Astronomical Institute, Moscow, informs us

that the designation KL Dra has been given to this dwarf nova, originally announced as SN 1998di (cf. IAUC 6982, 6983).

2000 March 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green